OpenStudy (anonymous):

fill in the answers! x + 8 - blank = 12 - blank x + 8 =12 THEN x =

OpenStudy (saifoo.khan):

x + 8 - blank = 12 - blank blank = 0 x+8 = 12 x = 12 - 8 x = 4

OpenStudy (anonymous):

blank is 0

OpenStudy (saifoo.khan):

Like i said.

OpenStudy (anonymous):

Thanks!!! there are 11 more of these!!!!

OpenStudy (anonymous):

LOL